# Simulation-based inference using surjective sequential neural likelihood estimation

### v2

@@ 4.3 Solar dynamo
4.4 Neural mass model
We also evaluate SSNL on the stochastic version of the Jansen-Rit neural mass model (Buckwar et al., 2020) which describes the collective electrical activity of a neurons by modelling interactions of cells. Neural mass models are highly used in practice and have contributed significantly to the understanding of the dynamics of populations of neurons (see Ableidinger et al. (2017); Rodrigues et al. (2021) for details). The model is a $6$-dimensional SDE of the form
$\begin{split}\mathrm{d}\begin{pmatrix}R_{t}\\ S_{t}\end{pmatrix}=&\begin{pmatrix}S_{t}\\ -\Gamma^{2}R_{t}-2\Gamma S_{t}+G_{\theta}(R_{t})\end{pmatrix}\mathrm{d}t\\ &+\begin{pmatrix}0\\ \Sigma_{\theta}\end{pmatrix}\mathrm{d}W_{t}\end{split}$
where $R=[Y_{1},Y_{2},Y_{3}]^{T}$, $S=[Y_{4},Y_{5},Y_{6}]^{T}$, $W_{t}$ is a Wiener process, $\Sigma_{\theta}$ is a diagonal covariance matrix, $G_{\theta}$ is a displacement vector, $\Gamma$ is a matrix, and $\theta$ is a four-dimensional random vector (see Appendix E.9 for details). For this experiment, SNPE-C on average outperforms SSNL (Figure 5 right column). However, with increasing sample sizes SSNL achieves as similar performance as SNPE-C. The inferences of SNL are significantly worse than for SSNL (more experimental results, including posterior distributions, can be found in Appendix F).
Figure 5: Solar dynamo and neural mass model evaluation (we show the MSE w.r.t. the prior sample $\theta_{0}$ that was used to simulate the observation $y_{0}$. Values are normalized by the minimum MSE in the entire data set).
@@ 5 Related work
Neural simulation-based inference is an emergent field with significant advances in the recent past. In addition to (sequential) neural approximations to the posterior, likelihood or likelihood-ratio, recent approaches have utilized flow matching and score-based models for posterior inference (Simons et al., 2023; Schmitt et al., 2023; Geffner et al., 2023), albeit mainly in a non-sequential manner. Score-based NPE methods are a fruitful research direction which replace the strict restriction on neural architectures of NFs. More recently, Gruner et al. (2023) proposed a new method that is targeted at models in which the posterior is conditioned on multiple observations simultaneously. Jia (2024) introduce a new family of methods, called neural quantile estimation, which uses quantile regression to either approximate the intractable posterior or likelihood functions of a model. Glaser et al. (2022); Pacchiardi & Dutta (2022) propose approaches to SBI using energy-based models as density estimators. Finally, Yao et al. (2023) propose a method based on stacking to combine the results of multiple posterior inferences, i.e., when multiple posterior approximations from different methods are available.
Loosely related to our work, Alsing & Wandelt (2018); Chen et al. (2021, 2023) have developed methods for SBI to compute summary statistics which, however, requires learning an additional embedding network while our approach learns an embedding and computes likelihood approximations in a single step and without computational overhead. Radev et al. (2023) discuss an approach that learns three networks for posterior approximations, likelihood approximations and summary statistic computation. Beck et al. (2022) discuss marginalization of data dimensions to see the influence of different covariates on the posterior distribution and propose an approach to find informative data dimensions.
In order to evaluate inferences of non-sequential procedures of models for which samples from the true posterior are not available, Linhart et al. (2023) have proposed a local procedure based on classifier tests. Similarly, Yao & Domke (2023) proposed a diagnostic related to, and with higher power than, simulation-based calibration (Talts et al. (2018); unfortunately, for neural likelihood methods we found them computationally infeasible to use, since they require either computing a vast number of permutation tests of which each requires learning a classifier or repeatedly sampling from the surrogate posterior for every model).
@@ 6 Conclusion
We introduced Surjective Sequential Neural Likelihood estimation, a new neural likelihood method in the family of methods for SBI. SSNL uses dimensionality-reducing surjections to embed the data in a lower-dimensional space while simultaneously learning the likelihood function. Especially when simulators produce high-dimensional data samples that can be embedded in a lower-dimensional space, we showed that our method performs favourably in contrast to several state-of-the-art baselines.
Reducing the dimensionality of the data has, however, other benefits: empirically we found that SSNL converges faster during training than SNL and that it was faster to draw posterior samples (since evaluating the likelihood in gradient-based methods such as HMC (Betancourt, 2017) is cheaper).
We anticipate that our method will complement other approaches to SBI well, but do not argue that it should be generally preferred over others. For instance, neural posterior methods, such as SNPE-C, often have excellent off-the-shelf performance. Similarly, neural ratio methods, such as SNRE-C, are typically easier to fit as flow-based methods, since they only require training a classifier. Consequently, the method of choice depends on the problem at hand.
Future research could elucidate how deeper (stacked) surjective architectures impact posterior inference and how, in general, more rigorous model selection procedures for neural SBI methods can be conducted. Building on the work by, e.g., Chen et al. (2021); Hoffmann & Onnela (2022), it would be interesting to research if likelihood estimation with dimensionality-reducing approaches can be improved via information-theoretic criteria, e.g., in order to train NNs more efficiently and ultimately to reduce the number of data simulations.
Societal impact
We do not anticipate any negative societal or ethical impacts of our work.
@@ A.1 Neural posterior estimation
Neural posterior estimation (SNPE) methods (Papamakarios & Murray, 2016; Lueckmann et al., 2017; Greenberg et al., 2019; Deistler et al., 2022; Wildberger et al., 2023) use a normalizing flow to directly target the posterior distribution thereby approximating $q_{f}(\theta|{y})\approx p(\theta|{y})$. SNPE-C (Greenberg et al., 2019)) uses the same sequential training procedure as SNL. In the first round, however, it optimizes the maximum likelihood objective $\mathbb{E}_{\mathcal{D}}\left[q_{f}(\theta|{y})\right]$ in each round. Subsequent rounds proceed by first composing a proposal prior as $\hat{p}^{r}(\theta)=q_{f}(\theta|{y}_{0})$, simulating new pairs $\{{y}_{n},\theta_{n}\}^{r}_{1\dots N}$ where $\theta_{n}\sim\hat{p}^{r}(\theta)$ and then re-training the NF. Since the parameters are sampled from the proposal prior $\hat{p}^{r}(\theta)$, the surrogate posterior would no longer target the true posterior $p(\theta|{y})$ but rather
@@
Greenberg et al. (2019) overcome this by deriving the new objective $\mathbb{E}_{\mathcal{D}}\left[\frac{1}{Z}q_{f}(\theta|{y})\frac{\hat{p}^{r}(\theta)}{p(\theta)}\right]$ which however requires the computation of a normalization constant $Z$.
SNPE-C can simulate posterior realizations by sampling from the normalizing flow base distribution first, and then propagating the samples through the flow layers. This can lead to posteriors that are outside of the prior bounds which need to be rejected and the procedure repeated until a sample of desired size has been taken. Specifically, if the prior distributions are constrained, e.g., containing scale parameters, or are very narrow, APT is known to exert ’leakage’, i.e., the posterior approximation might produce samples that are not within the prior bounds. In this case, the rejection rate of posterior samples is elevated, for instance, as reported in Durkan et al. (2020) or Glöckler et al. (2022), the latter of which having observed rejection rates of up to $99\%$, which necessitates the use of MCMC methods instead. Leakage significantly reduces the usefulness of SNPE methods in comparison to SNL where draws are generated using MCMC in the first place. Furthermore, for structured data sets, e.g., time series data, SNPE-C requires facilitating a second neural network to embed the data before conditioning which increases the number of effective parameters.
@@ A.2 Likelihood ratio estimation
Neural likelihood ratio estimation (NRE) methods (Hermans et al., 2020; Durkan et al., 2020; Miller et al., 2022; Delaunoy et al., 2022)) learn the likelihood-to-evidence ratio ${r}({y},\theta)=\frac{p({y}|\theta)}{p({y})}=\frac{p(\theta|y)}{p(\theta)}$ and then build a surrogate posterior $\hat{p}(\theta|{y})=\hat{r}({y},\theta)p(\theta)$. A major advantage of NRE methods is, that they do not need to train a model that estimates a density using normalizing flow which often brings significant computational and numerical advantages. While NRE-C (Miller et al., 2022) has been proposed in a non-sequential scenario, it is also possible to derive posterior distributions sequentially (SNRE-C; see, e.g., Tejero-Cantero et al. (2020)). In this case, since a proposal posterior $p^{r}(\theta|y_{0})$ is derived after round $r$, which changes the joint distribution to $p(y|\theta)p^{r}(\theta|y_{0})$, the estimated ratio becomes
@@ A.3 Notes
The idea of using non-trivial embedding networks, such as CNNs or LSTMs for NPE and NRE methods is not new (see e.g., Greenberg et al. (2019) or the notebooks of the SBI Python package [^3] https://sbi-dev.github.io/sbi/tutorial/05_embedding_net/). This requires an additional neural network and consequently increases the number of total parameters. We, on the other hand, do dimensionality reduction and likelihood estimation in one step and with one network.
@@ Appendix B Mathematical derivations
The derivation of the surjection layer used in SSNL largely follows the SurVAE framework of Nielsen et al. (2020) and Klein et al. (2021). The SurVAE framework models the log-probability $\log p({y})$ of a $P$-dimensional data point ${y}\in\mathcal{Y}$ as
@@
Nielsen et al. (2020) define the likelihood contribution for inference surjections as
@@
By observing (see also Appendix A of Nielsen et al. (2020) and main manuscript Klein et al. (2021)) that the composition of a differentiable function $g$ with a Dirac $\delta$ function and a bijection $f$ is
@@
(a) Bijection.
(b) Surjection.
Figure 6: Comparison between the $k$th masked autoregressive bijection and masked autoregressive surjection layers $f_{k}$, respectively (green: conditioning subvector; blue: value that is transformed, violet: additional conditioning subvector used by surjection layer, red: conditioner and transformer functions). (a) The $k$ masked autoregressive flow requires a conditioner $c_{k}$, a transformer $\tau$ and a conditioning variable $\theta$. (b) A surjection computes the same operation on a subvector $z_{k,+}$. It, in addition uses the subvector $z_{k,-}$ for conditioning.
@@ Appendix C Implementation details
Surjection layers can be implemented in a straight-forward manner by extending the bijection layers of conventional machine libraries. Below, we demonstrate the implementation of a conditional affine masked autoregressive surjective flow that uses an affine MAF (Papamakarios et al., 2017), called AffineMaskedAutoregressive as a super class.
@@
where MADE is a masked autoencoder for density estimation (Germain et al., 2015), decoder corresponds to the conditional density $p(y_{-}|z,\theta)$.
@@ D.1 Implementation details
For SSNL and SNL, we used custom implementations using the JAX-based libraries Haiku, Distrax, TensorFlow Probability and the optimization framework Optax (Bradbury et al., 2018; Babuschkin et al., 2020; Dillon et al., 2017). For sampling, SSNL and SNL use the JAX-based library BlackJAX (Cabezas et al., 2023). For SNPE-C and SNRE-C, we used the SBI toolbox (Tejero-Cantero et al., 2020) which offers high-quality implementations of SBI methods and model diagnostics. We simulate data from stochatic differential equations using the JAX-based package Diffrax (Kidger, 2021).
@@ D.2 Training and sampling
We trained each model using an Adam optimizer with fixed learning rate of $r=0.0001$ and momentums $b_{1}=0.9$ and $b_{2}=0.999$. Each experiment uses a mini-batch size of $100$. The optimizer is run until a maximum of $2000$ epochs is reached or no improvement on a validation set can be observed for $10$ consecutive iterations. The validation set consists of $10\%$ of the entire data set, while the other $90\%$ are used for training. For each round, we start training the neural network from scratch and do not continue from the previously learned state.
@@
For SSNL and SNL, we used the No-U-turn sampler (Hoffman et al., 2014) from the sampling library BlackJAX to sample from the intermediate and final posterior distributions $p^{r}(\theta|{y}_{0})$ using $4$ chains of a fixed length of 10 0001000010\,00010 000 each of which the first $5000$ iterations are discarded as burn-in per chain. For SNPE-C and SNRE-C experiments, we use the slice sampler of the SBI toolbox for sampling (which we do in lieu of rejection sampling to avoid leakage; see the Appendix A).Samples from the "true" posterior distribution have been drawn using TensorFlow Probability’s slice sampler where we used $10$ chains of length 20 0002000020\,00020 000 of which we discarded the first 10 0001000010\,00010 000 as burn-in. Convergence of the true posteriors in this case has been diagnosed using the potential scale reduction factor (Gelman & Rubin, 1992; Vehtari et al., 2021), effective sample size calculations, and conventional graphical diagnostics (Gabry et al., 2019).
@@ D.3 Neural network architectures
For all experiments and evaluated methods, we used the same neural network architectures. We followed the neural network architectures as described in Greenberg et al. (2019); Papamakarios et al. (2019); Miller et al. (2022) and tried to keep the number of total parameters of each model as comparable as possible to allow for an unbiased evaluation.
@@
The SSNL architectures use a total of $K=5$ layers, the third of which is a surjection layer with reduction factors of $25\%$, $50\%$ or $75\%$ which we chose arbitrarily. For instance, for a reduction factor of $25\%$, we take the initial dimensionality $P$ and reduce it to $Q=\lfloor 0.25*P\rfloor$. Each of the layers is parameterized by a MAF which uses a MADE network with two layers with $50$ neurons each as conditioner (Germain et al., 2015; Papamakarios et al., 2017). The MAFs use tanh activation functions. The conditional densities are parameterized using a two-layer MLP with tanh activation functions. Between each MAF layer, we add a permutation layer that reverses the vector dimensions. In practice, we assume that optimizing the number of surjection layers and their reduction factors is advisable. This can, for instance, be done empirically by examining the likelihood profiles during training or by reducing to the same order of magnitude as required summary statistics.
@@
SNL uses a total of $K=5$ layers. Each of the layers is parameterized by a MAF which uses a MADE network with two layers with $50$ neurons each as conditioner (Germain et al., 2015; Papamakarios et al., 2017) with permutations in between which reverse the vector dimensions. SNL uses tanh activation functions.

### v3

@@ 5 Related work
Neural simulation-based inference is an emergent field with significant advances in the recent past. In addition to (sequential) neural approximations to the posterior (Papamakarios and Murray,, 2016; Lueckmann et al.,, 2017; Greenberg et al.,, 2019; Deistler et al.,, 2022; Wildberger et al.,, 2023), likelihood (Papamakarios et al.,, 2019; Glöckler et al.,, 2022) or likelihood-ratio (Cranmer et al.,, 2015; Durkan et al.,, 2020; Hermans et al.,, 2020; Thomas et al.,, 2022; Delaunoy et al.,, 2022; Miller et al.,, 2022), recent approaches have utilized flow matching and score-based models for posterior inference (Schmitt et al.,, 2024; Geffner et al.,, 2023; Wildberger et al.,, 2023), albeit mainly in a non-sequential manner. Score-based NPE methods are a fruitful research direction which do not restrict the architecture of the neural network architecture. More recently, Gruner et al., (2023) proposed a new method that is targeted at models in which the posterior is conditioned on multiple observations simultaneously. Jia, (2024) introduce a new family of methods, called neural quantile estimation, which uses quantile regression to either approximate the intractable posterior or likelihood functions of a model. Glaser et al., (2022); Pacchiardi and Dutta, (2022) propose approaches to SBI using energy-based models as density estimators. Finally, Yao et al., (2024) propose a method based on stacking to combine the results of multiple posterior inferences, i.e., when multiple posterior approximations from different methods are available.
Related to our work, Alsing and Wandelt, (2018); Chen et al., (2021, 2023) have developed methods for SBI to compute summary statistics which, however, requires learning an additional embedding network while our approach learns an embedding and computes likelihood approximations in a single step and without computational overhead. Radev et al., (2023) discuss an approach that learns three networks for posterior approximations, likelihood approximations and summary statistic computation. Beck et al., (2022) discuss marginalization of data dimensions to see the influence of different covariates on the posterior distribution and propose an approach to find informative data dimensions.
In order to evaluate inferences of non-sequential procedures of models for which samples from the true posterior are not available, Linhart et al., (2023) have proposed a local procedure based on classifier tests. Similarly, Yao and Domke, (2023) proposed a diagnostic related to, and with higher power than, simulation-based calibration (Talts et al., (2018); unfortunately, for neural likelihood methods we found them computationally infeasible to use, since they require either computing a vast number of permutation tests of which each requires learning a classifier or repeatedly sampling from the surrogate posterior for every model).
@@ 6 Conclusion
We introduced Surjective Sequential Neural Likelihood estimation, a new neural likelihood method for simulation-based inference for high-dimensional data. SSNL uses dimensionality-reducing surjections to embed the data in a lower-dimensional space while simultaneously learning the likelihood function.
SSNL performs particularly well when applied to high-dimensional time series data outperforming state-of-the-art methods and is on par with them in other experiments, making it an excellent off-the-shelf estimator for high-dimensional data sets. As a limitation, we identified that due to introducing an additional conditional density, SSNL in some cases requires a higher sample size than other flow-based models like SNL or SNPE-C.
We anticipate that our method will complement other approaches in SBI well, but do not argue that it should be generally preferred over others. For instance, when posteriors have a simple geometry and the likelihood function is complicated to approximate, posterior estimators like SNPE-C often provide superior performance. Similarly, neural ratio methods, such as SNRE-C, are typically easier to fit as flow-based methods, since they only require training a classifier.
Future research could elucidate how deeper (stacked) surjective architectures impact posterior inference and if more suitable surjective architectures can be developed for SBI. The application of information-theoretic criteria to more rigorously determine the dimensionality of the latent space, for instance, as done in Chen et al., (2023), is another possible research direction.
Acknowledgements.
This research was supported by the Swiss National Science Foundation (Grant No. $200021\_208249$). We thank all anonymous reviewers for their feedback which helped to improve the manuscript.
@@ A.1 Neural posterior estimation
Neural posterior estimation (SNPE) methods (Papamakarios and Murray,, 2016; Lueckmann et al.,, 2017; Greenberg et al.,, 2019; Deistler et al.,, 2022; Wildberger et al.,, 2023) use a normalizing flow to directly target the posterior distribution thereby approximating $q_{f}(\theta|{y})\approx p(\theta|{y})$. SNPE-C (Greenberg et al.,, 2019)) uses the same sequential training procedure as SNL. In the first round, however, it optimizes the maximum likelihood objective $\mathbb{E}_{\mathcal{D}}\left[q_{f}(\theta|{y})\right]$ in each round. Subsequent rounds proceed by first composing a proposal prior as $\hat{p}^{r}(\theta)=q_{f}(\theta|{y}_{0})$, simulating new pairs $\{{y}_{n},\theta_{n}\}^{r}_{1\dots N}$ where $\theta_{n}\sim\hat{p}^{r}(\theta)$ and then re-training the NF. Since the parameters are sampled from the proposal prior $\hat{p}^{r}(\theta)$, the surrogate posterior would no longer target the true posterior $p(\theta|{y})$ but rather
@@
Greenberg et al., (2019) overcome this by deriving the new objective $\mathbb{E}_{\mathcal{D}}\left[\frac{1}{Z}q_{f}(\theta|{y})\frac{\hat{p}^{r}(\theta)}{p(\theta)}\right]$ which however requires the computation of a normalization constant $Z$.
SNPE-C can simulate posterior realizations by sampling from the normalizing flow base distribution first, and then propagating the samples through the flow layers. This can lead to posteriors that are outside of the prior bounds which need to be rejected and the procedure repeated until a sample of desired size has been taken. Specifically, if the prior distributions are constrained, e.g., containing scale parameters, or are very narrow, APT is known to exert ’leakage’, i.e., the posterior approximation might produce samples that are not within the prior bounds. In this case, the rejection rate of posterior samples is elevated, for instance, as reported in Durkan et al., (2020) or Glöckler et al., (2022), the latter of which having observed rejection rates of up to $99\%$, which necessitates the use of MCMC methods instead. Leakage significantly reduces the usefulness of SNPE methods in comparison to SNL where draws are generated using MCMC in the first place. Furthermore, for structured data sets, e.g., time series data, SNPE-C requires facilitating a second neural network to embed the data before conditioning which increases the number of effective parameters.
@@ A.2 Likelihood ratio estimation
Neural likelihood ratio estimation (NRE) methods (Hermans et al.,, 2020; Durkan et al.,, 2020; Miller et al.,, 2022; Delaunoy et al.,, 2022)) learn the likelihood-to-evidence ratio ${r}({y},\theta)=\frac{p({y}|\theta)}{p({y})}=\frac{p(\theta|y)}{p(\theta)}$ and then build a surrogate posterior $\hat{p}(\theta|{y})=\hat{r}({y},\theta)p(\theta)$. A major advantage of NRE methods is, that they do not need to train a model that estimates a density using normalizing flow which often brings significant computational and numerical advantages. While NRE-C (Miller et al.,, 2022) has been proposed in a non-sequential scenario, it is also possible to derive posterior distributions sequentially (SNRE-C; see, e.g., Tejero-Cantero et al., (2020)). In this case, since a proposal posterior $p^{r}(\theta|y_{0})$ is derived after round $r$, which changes the joint distribution to $p(y|\theta)p^{r}(\theta|y_{0})$, the estimated ratio becomes
@@ A.3 Notes
The idea of using non-trivial embedding networks, such as CNNs or LSTMs for NPE and NRE methods is not new (see e.g., Greenberg et al., (2019) or the notebooks of the SBI Python package [^2] https://sbi-dev.github.io/sbi/tutorial/05_embedding_net/). This requires an additional neural network and consequently increases the number of total parameters. We, on the other hand, do dimensionality reduction and likelihood estimation in one step and with one network.
@@ Appendix B Mathematical derivations
The derivation of the surjection layer used in SSNL largely follows the SurVAE framework of Nielsen et al., (2020) and Klein et al., (2021). The SurVAE framework models the log-probability $\log p({y})$ of a $P$-dimensional data point ${y}\in\mathcal{Y}$ as
@@
Nielsen et al., (2020) define the likelihood contribution for inference surjections as
@@
By observing (see also Appendix A of Nielsen et al., (2020) and main manuscript Klein et al., (2021)) that the composition of a differentiable function $g$ with a Dirac $\delta$ function and a bijection $f$ is
@@ Appendix C Implementation details
Surjection layers can be implemented in a straight-forward manner by extending the bijection layers of conventional machine libraries. Below, we demonstrate the implementation of a conditional affine masked autoregressive surjective flow that uses an affine MAF (Papamakarios et al.,, 2017), called AffineMaskedAutoregressive as a super class.
@@
where MADE is a masked autoencoder for density estimation (Germain et al.,, 2015), decoder corresponds to the conditional density $p(y_{-}|z,\theta)$.
@@ D.1 Implementation details
All models are implemented using the Python packages sbijax, (Dirmeier et al.,, 2024), surjectors (Dirmeier,, 2024), the SBI toolbox (Tejero-Cantero et al.,, 2020), and the Deepmind JAX ecosystem (Bradbury et al.,, 2018; Babuschkin et al.,, 2020). We simulate data from stochastic differential equations using the package Diffrax (Kidger,, 2021).
@@ D.2 Training and sampling
We trained each model using an Adam optimizer with fixed learning rate of $r=0.0001$ and momentums $b_{1}=0.9$ and $b_{2}=0.999$. Each experiment uses a mini-batch size of $100$. The optimizer is run until a maximum of $2000$ epochs is reached or no improvement on a validation set can be observed for $10$ consecutive iterations. The validation set consists of $10\%$ of the entire data set, while the other $90\%$ are used for training. For each round, we start training the neural network from scratch and do not continue from the previously learned state. Each model was trained on a HPC computing cluster using a single node consisting of two 18 core Broadwell CPUs (Intel Xeon E5-2695 v4).
@@
For SSNL and SNL, we used the No-U-turn sampler (Hoffman and Gelman,, 2014) from the sampling library BlackJAX to sample from the intermediate and final posterior distributions $p^{r}(\theta|{y}_{0})$ using $4$ chains of a fixed length of 10 0001000010\,00010 000 each of which the first $5000$ iterations are discarded as burn-in per chain. For SNPE-C and SNRE-C experiments, we use the slice sampler of the SBI toolbox for sampling (which we do in lieu of rejection sampling to avoid leakage; see the Appendix A).Samples from the "true" posterior distribution have been drawn using TensorFlow Probability’s slice sampler where we used $10$ chains of length 20 0002000020\,00020 000 of which we discarded the first 10 0001000010\,00010 000 as burn-in. Convergence of the true posteriors in this case has been diagnosed using the potential scale reduction factor (Gelman and Rubin,, 1992; Vehtari et al.,, 2021), effective sample size calculations, and conventional graphical diagnostics (Gabry et al.,, 2019).
@@ D.3 Neural network architectures
For all experiments and evaluated methods, we used the same neural network architectures. We followed the neural network architectures as described in Greenberg et al., (2019); Papamakarios et al., (2019); Miller et al., (2022) and tried to keep the number of total parameters of each model as comparable as possible to allow for an unbiased evaluation.
@@
The SSNL architectures use a total of $K=5$ layers, the third of which is a surjection layer with reduction factors of $25\%$, $50\%$ or $75\%$ which we chose arbitrarily. For instance, for a reduction factor of $25\%$, we take the initial dimensionality $P$ and reduce it to $Q=\lfloor 0.25*P\rfloor$. Each of the layers is parameterized by a MAF which uses a MADE network with two layers with $50$ neurons each as conditioner (Germain et al.,, 2015; Papamakarios et al.,, 2017). The MAFs use tanh activation functions. The conditional densities are parameterized using a two-layer MLP with tanh activation functions. Between each MAF layer, we add a permutation layer that reverses the vector dimensions. In practice, we assume that optimizing the number of surjection layers and their reduction factors is advisable. This can, for instance, be done empirically by examining the likelihood profiles during training or by reducing to the same order of magnitude as required summary statistics.
@@
SNL uses a total of $K=5$ layers. Each of the layers is parameterized by a MAF which uses a MADE network with two layers with $50$ neurons each as conditioner (Germain et al.,, 2015; Papamakarios et al.,, 2017) with permutations in between which reverse the vector dimensions. SNL uses tanh activation functions.
@@
SNASS
To keep the number of parameters as equal as possible, SNASS uses a normalizing flow using three flow layers each consisting of a MADE with two layers and 50 nodes each. SNASS uses as summary and critic networks two MLPs with two hidden layers and 50 nodes each. All activation functions are tanhs.
SNASSS
Similarly, SNASS uses a normalizing flow using two flow layers each consisting of a MADE with two layers and 50 nodes each. SNASSS uses as summary and critic networks three MLPs with a single hidden layer and 50 nodes. All activation functions are tanhs.